How many liters of a 30% acid solution should be mixed with 5 liters of a 21% acid solution to produce a 27% solution?
Answer by checkley77(12844)

.30x+.21*5=.27(x+5)
.30x+1.05-.27x+1.35
.03x=1.35-1.05
.03x=.30
x=.30/.03
x=10 liters of 30% solution is used.
Proof:
.30*10+.21*5=.27(10+5)
3+1.05=.27*15
4.05=4.05